Problem 9-3A On January 1, 2019, Blue Spruce Company purchased the following two machines for use in its production process Machine A: The cash price of this machine was $56,500 Related expenditures included: sales tax S2 55 , shipping costs S10 insurance during shipping s installation and testing costs $90, and $200 of oil and lubricants to be used with the machinery during its first year of operations. Blue Spruce estimates that the useful life of the machine is 5 years with a $4,350 salvage value remaining at the end of that time period. Assume that the straight-line method of depreciation is used Machine B: The recorded cost of this machine was $180,000. Blue Spruce estimates that he use ul life of the machines·ers with a SL at the end of that time period. LO sa age value remaining Prepare the following for Machine A. (Round answers to O decimal places, e.g. 2,125. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter O for the amounts.) 1. The journal entry to record its purchase on January 1, 2019. 2. The journal entry to record annual depreciation at December 31, 2019 No. Account Titles and Explanation 1. Equipment Debit 59490 Cash 9490 2. Depreciation Expense Accumulated Depreciation-Equipment
Answer #1

Depreciation as per straight Line Method- Cost of Asset - Salvage Value at the end of useful life

Life of Asset

Depreciation as per Straight Line Method (Machine A) - (59490 - 4350) / 5

= 11028.

Depreciation as per straight Line Method (Machine B) 2019, 2020, 2021, 2022 -

= (180000 - 10000) / 4 = 42500

Depreciation as per declining balance Method 2019- 190000 * 50% = 95000

Depreciation as per declining balance Method 2020- 95000 * 50% = 47500

Depreciation as per declining balance Method 2021- 47500 * 50% = 23750

Depreciation as per declining balance Method 2022- 23750 * 50% = 11875

Depreciation as per units of activity method 2019 - 190000 * 45500 / 110100 = 78519.52

Depreciation as per units of activity method 2020 - 190000 * 32500 / 110100 = 56085.37

Depreciation as per units of activity method 2021 - 190000 * 19500 / 110100 = 33651.22

Depreciation as per units of activity method 2019 - 190000 * 12600 / 110100 = 21743.86

Highest Depreciation in the 1st Year (2019) on Machine B - Declining Balance Method (95000)

Highest Depreciation in the 4th Year (2022) on Machine B - Straight Line Method (42500)

Highest total Depreciation over the 4 Year period on Machine B - Units of Activity Method (190000)